ABC Company just paid annual dividend of $5 today. The dividend is expected to grow at 3% for the next 5 years, then it will grow at 5% in perpetuity. If stocks of similar company earn 9% annual return, what is the price of a share of ABC Company stock?
Answer #1

Particulars Dividend Dividend Growth Rate Year Ol Year 1 $5.00 $5.15 3% Year 21 Year 3 $5.30 $5.46 3% 3% Year 4 $5.63 3% Year

Particulars Dividend Dividend Growth Rate =C2+1 =C3*(1+D4) =D2+1 =D3*(1+E4) 0.03 =E2+1 =E3*(1+F4) 0.03 =F2+1 =F3*(1+G4) 0.03